# The Riemann Hypotheses is false

### v26

@@
Introduction. The Riemann zeta function is a function of the complex variable $s$, defined in the half-plane $\Re(s)>1$ by $\zeta(s):=\sum_{n=1}^{\infty}n^{-s}$ and in the whole complex plane by holomorphic continuation. Euler noticed that for $\Re(s)>1$, $\zeta(s)$ can be expressed as a product $\prod_{p}(1-p^{-s})^{-1}$ over the entire set of primes, which entails that $\zeta(s)\neq 0$ for $\Re(s)>1$. It can be shown that $\zeta(s)$ extends to $\mathbb{C}$ as a meromorphic function with only a simple pole at $s=1$, with residue $1$, and satisfies the functional equation $\xi(s)=\xi(1-s)$, where $\xi(z)=\frac{1}{2}z(z-1)\pi^{-z/2}\Gamma(\frac{1}{2}z)\zeta(z)$ and $\Gamma(w)=\int_{0}^{\infty}e^{-x}x^{w-1}\mathrm{d}x$. Let $\chi$ be a primitive Dirichlet character and $L(s,\chi)$ be the corresponding L-function. Most of the analytic properties for $\zeta$ are also known to hold for $L(s,\chi)$ [5, Chapter 10]. Let $\Theta_{\chi}$ be the supremum of the real parts of the zeros of $L(s,\chi)$. The Generalised Riemann Hypothesis (GRH) is equivalent to the statement that $\Theta_{\chi}=\frac{1}{2}$ for each $\chi$, and the Generalised Prime Number Theorem is equivalent to the statement that $L(s,\chi)\neq 0$ at $\Re(s)=1$ for each $\chi$. For a far more thorough discussion of the RH and GRH, the interested reader is kindly referred to Borwein et al [1].
Definitions. Let: $\rho$ denote a non-trivial zero of the Riemann zeta function, $\gamma=\Im(\rho)$ and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s. Denote by $\Lambda$ the von Mangoldt function.
@@
Suppose that $\Theta<1$ and fix $0<\varepsilon<1-\Theta$. Let $T\in\mathbb{R}_{>0}$ be arbitrarily large and $X\geq 2$. By Lemma 12.2 of [5], we have $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX)\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$. Thus we can choose $T$ and $H\in[3T,3T+1]$ such that
@@
$\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{\Theta+\varepsilon+2iT}^{\Theta+\varepsilon+iH}+\int_{\Theta+\varepsilon+iH}^{-1+iH}+\int_{-1+iH}^{-1+2iT}+\int_{-1+2iT}^{\Theta+\varepsilon+2iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
@@
say. From now on, let $2T\leq t\leq H$ and $k=k(\varepsilon)>6$ be a fixed real number such that $k(1-\Theta-\varepsilon)\geq 3$. For $0<a<b\leq 2a$, note that
@@
$\sum_{\frac{T}{\log^{k}T}<n\leq T}n^{it}$ $=\Big{(}\sum_{\frac{T}{\log^{k}T}<n\leq\frac{2T}{\log^{k}T}}+\sum_{\frac{2T}{\log^{k}T}<n\leq\frac{4T}{\log^{k}T}}+\cdots+\sum_{\frac{T}{2}<n\leq T}\Big{)}n^{it}$ (5)
$\ll T^{\frac{1}{2}}\log^{k+1}T.$ (6)
By partial summation, one deduces from (6) that
$\sum_{\frac{T}{\log^{k}T}<n\leq T}n^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{k+1}T$ (7)
@@
for $\sigma\in\mathbb{R},|\sigma|\leq A$ where $A>0$ is a constant. Combining (1) and (7) yields $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{2}+I_{4}\ll T^{1/2}.$ Let $\sigma\geq\varepsilon/2$. By the log-derivative of the functional equation for $\zeta$ (see e.g. [3, p.403]), note that
$\frac{\zeta^{\prime}}{\zeta}(-\sigma+it)\ll\log t.$ (8)
Combining (7) and (8) gives $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{3}\ll T^{1/2}\log^{k+2}T.$ Thus we arrive at
$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{k+2}T.$ (9)
Let $s=\Theta+\varepsilon+iY$ and $c=1-\Theta-\varepsilon+\frac{1}{\log T}>0$, where $Y\in[2T,H]$. By Theorem 3.12 of [6], note that
$-\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\Lambda(m)}{m^{s}}=\frac{1}{2\pi i}\int_{c-4iT}^{c+4iT}\frac{\zeta^{\prime}}{\zeta}(s+w)\frac{(T/2\log^{k}T)^{w}}{w}\mathrm{d}w+O(T^{c-1}\log^{1-kc}T).$ (10)
Notice that the above integrand has simple poles at $w=0,w=\rho$ and $w=1-s$. Hence, shifting the line of integration in (10) to $\Re(w)=-1$ then invoking (1) and (8) to bound the integrals from Cauchy’s residue theorem, gives
$\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\Lambda(m)}{m^{s}}-\frac{(T/2\log^{k}T)^{1-s}}{1-s}+\sum_{2T\leq\gamma_{T}\leq H}\frac{(T/2\log^{k}T)^{\rho-s}}{\rho-s}+O(T^{c-1}\log^{1-kc}T).$ (11)
Putting (11) into $I_{1}$ gives
$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m\leq\frac{T}{2\log^{k}T}}\Bigg{(}\frac{n^{\Theta+\varepsilon-1+2iT}\Lambda(m)}{m^{\Theta+\varepsilon+2iT}\log(n/m)}-\frac{n^{\Theta+\varepsilon-1+iH}\Lambda(m)}{m^{\Theta+\varepsilon+iH}\log(n/m)}\Bigg{)}$ (12)
$-\frac{1}{2\pi i}\int_{\Theta+\varepsilon+2iT}^{\Theta+\varepsilon+iH}\frac{(T/2n\log^{k}T)^{1-s}}{1-s}\mathrm{d}s+\frac{1}{2\pi i}\sum_{2T\leq\gamma_{T}\leq H}(T/2\log^{k}T)^{\rho_{T}}\int_{\Theta+\varepsilon+2iT}^{\Theta+\varepsilon+iH}\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s$ (13)
$+O(n^{\Theta+\varepsilon-1}T^{c}\log^{2-kc}T).$ (14)
By the residue theorem again, we have
$\frac{1}{2\pi i}\int_{\Theta+\varepsilon+2iT}^{\Theta+\varepsilon+iH}\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s$ $=\sum_{2T\leq\gamma_{T}\leq H}(T/2\log^{k}T)^{-\rho_{T}}n^{\rho_{T}-1}-\frac{1}{2\pi i}\Bigg{(}\int_{\Theta+\varepsilon+iH}^{\Theta+\varepsilon+2iT}+\int_{\Theta+\varepsilon+2iT}^{-\varepsilon+2iT}+$ (15)
$\int_{-\varepsilon+2iT}^{-\varepsilon+iH}\Bigg{)}\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s$ (16)
$=\sum_{2T\leq\gamma_{T}\leq H}(T/2\log^{k}T)^{-\rho_{T}}n^{\rho_{T}-1}-\frac{1}{2\pi i}\Bigg{(}\int_{\Theta+\varepsilon+iH}^{\Theta+\varepsilon+2iT}+\int_{\Theta+\varepsilon+2iT}^{-\varepsilon+2iT}\Bigg{)}$ (17)
$\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s+O(n^{-1-\varepsilon}T^{-\varepsilon}).$ (18)
Inserting (18), (17), (16) and (15) into the summation in (13), gives
$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m\leq\frac{T}{2\log^{k}T}}\Bigg{(}\frac{n^{\Theta+\varepsilon-1+2iT}\Lambda(m)}{m^{\Theta+\varepsilon+2iT}\log(n/m)}-\frac{n^{\Theta+\varepsilon-1+iH}\Lambda(m)}{m^{\Theta+\varepsilon+iH}\log(n/m)}\Bigg{)}$ (19)
$-\frac{1}{2\pi i}\int_{\Theta+\varepsilon+2iT}^{-\varepsilon+2iT}\frac{(T/2n\log^{k}T)^{1-s}}{1-s}\mathrm{d}s+\sum_{2T\leq\gamma_{T}\leq H}(n^{\rho_{T}-1}+O(n^{-1-\varepsilon}T^{-\varepsilon})).$ (20)
Since $1\leq m\leq\frac{T}{2\log^{k}T}$, note that $0<\frac{1}{\log(n/m)}\leq\frac{1}{\log 2}$ uniformly for $\frac{T}{\log^{k}T}\leq n\leq T$. Hence by partial summation, it follows from (6) that
$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{n^{2iT}}{m^{\Theta+\varepsilon+2iT}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{k+1}T}{m^{\Theta+\varepsilon}\log(T/m)}$ (21)
uniformly for $2\leq m\leq\frac{T}{2\log^{k}T}$. The same bound holds for $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{n^{iH}}{m^{\Theta+\varepsilon+iH}\log(n/m)}.$ By partial summation again, it follows from (7) that
$\sum_{\frac{T}{2\log^{k}T}\leq n\leq T}\int_{\Theta+\varepsilon+2iT}^{-\varepsilon+2iT}\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s\ll T^{1/2}\log^{k+2}T.$ (22)
@@
$\sum_{n\leq T}n^{\rho_{T}-1}=\zeta(1-\rho_{T})+\frac{T^{\rho_{T}}}{\rho_{T}}+O(T^{\Theta-1})\ll T^{\Theta-1}$ (23)
uniformly for $2T\leq\gamma_{T}\leq H$. Hence
$\sum_{1\leq n\leq T}\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=\sum_{2T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}\ll T^{\Theta}\log T.$ (24)
Recall that $H=3T+O(1)$ and $kc>3$. Summing (19) and (20) for $\frac{T}{\log^{k}T}\leq n\leq T$ and invoking the above estimates therefore yields
$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{1}$ $=-\frac{T}{2\pi}\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{\Lambda(n)}{n}+O\Bigg{(}T^{\frac{1}{2}}\log^{k+1}T\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\log m}{m^{\Theta+\varepsilon}\log(T/m)}\Bigg{)}+o(T)$ (25)
$=-\frac{T}{2\pi}\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{\Lambda(n)}{n}+o(T)$ (26)
$=-\frac{kT}{2\pi}\log\log T+o(T)$ (27)
since $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ [5, p.182], where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant. By Theorem 1 of [3], we have
$\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ (28)
uniformly for $2\leq n\ll T.$ Define $N(y):=\sum_{0<\gamma\leq y}1$. Note that $N(y)=(y/2\pi)\log(y/2\pi)-(y/2\pi)+O(\log y)$ [5, Corollary 14.3]. Thus from (28) we obtain
$\sum_{1\leq n\leq\frac{T}{\log^{k}T}}\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=N(H)-N(2T)-\frac{T}{2\pi}\sum_{2\leq n\leq\frac{T}{\log^{k}T}}\frac{\Lambda(n)}{n}+o(T)$ (29)
$=\frac{kT}{2\pi}\log\log T+\frac{T}{2\pi}(B-1+\log 2)+o(T),$ (30)
since $H=3T+O(1)$. Combining (30), (27) and (9) gives
$\sum_{1\leq n\leq T}\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\sum_{2T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}$ (31)
$=\Big{(}\sum_{2T\leq\gamma_{T}\leq H}\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{k}T}}+\sum_{\frac{T}{\log^{k}T}<n\leq T}\Big{)}\Big{)}n^{\rho_{T}-1}$ (32)
$\gg T.$ (33)
But (33) is a contradiction to (24), thus we deduce that our supposition must be false. ∎
@@
Remark 1. Notice that our argument actually shows that, for all large enough $T\in\mathbb{R}^{+}$, there exists some $\rho_{T}$ with $12T\leq\gamma_{T}\leq 3T+$ 1 and $\Re(\rho_{T})=1-o(1)$.
@@
(i) The functional equation for $\zeta$,
(ii) Theorem 3.12 of [6], and
(iii) Theorem 4.11 of [6].
Both (i) and (iii) do not hold for Beurling zeta functions corresponding to Beurling number systems whose integers are not the ordinary integers. Hence our approach wouldn’t work for such, whose analogues of the RH are known to be true. Note that both (i) and (ii) also do not hold for the Weil zeta functions, which are also known to satisfy the RH.

### v27

@@
 [^1] Data availability statement: My manuscript has no associated data. Introduction. The Riemann zeta function is a function of the complex variable $s$, defined in the half-plane $\Re(s)>1$ by $\zeta(s):=\sum_{n=1}^{\infty}n^{-s}$ and in the whole complex plane by holomorphic continuation. Euler noticed that for $\Re(s)>1$, $\zeta(s)$ can be expressed as a product $\prod_{p}(1-p^{-s})^{-1}$ over the entire set of primes, which entails that $\zeta(s)\neq 0$ for $\Re(s)>1$. It can be shown that $\zeta(s)$ extends to $\mathbb{C}$ as a meromorphic function with only a simple pole at $s=1$, with residue $1$, and satisfies the functional equation $\xi(s)=\xi(1-s)$, where $\xi(z)=\frac{1}{2}z(z-1)\pi^{-z/2}\Gamma(\frac{1}{2}z)\zeta(z)$ and $\Gamma(w)=\int_{0}^{\infty}e^{-x}x^{w-1}\mathrm{d}x$. Let $\chi$ be a primitive Dirichlet character and $L(s,\chi)$ be the corresponding L-function. Most of the analytic properties for $\zeta$ are also known to hold for $L(s,\chi)$ [5, Chapter 10]. Let $\Theta_{\chi}$ be the supremum of the real parts of the zeros of $L(s,\chi)$. The Generalised Riemann Hypothesis (GRH) is equivalent to the statement that $\Theta_{\chi}=\frac{1}{2}$ for each $\chi$, and the Generalised Prime Number Theorem is equivalent to the statement that $L(s,\chi)\neq 0$ at $\Re(s)=1$ for each $\chi$. For a far more thorough discussion of the RH and GRH, the interested reader is kindly referred to Borwein et al [1].
Definitions. Let: $\rho$ denote a non-trivial zero of $\zeta$, $\gamma=\Im(\rho)$ and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s.
@@
Suppose that $\Theta<1$ and fix an arbitrarily small $0<\varepsilon<1-\Theta$. Let $T\in\mathbb{R}_{>0}$ be arbitrarily large and $X\geq 2$. By Lemma 12.2 of [5], we have $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX)\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$. Thus we can choose $T$ and $H\in[2T,2T+1]$ such that
@@
$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{1+\frac{1}{\log n}+iT}^{1+\frac{1}{\log n}+iH}+\int_{1+\frac{1}{\log n}+iH}^{-1+iH}+\int_{-1+iH}^{-1+iT}+\int_{-1+iT}^{1+\frac{1}{\log n}+iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
@@
say. From now on, let $T\leq t\leq H$. For $0<a<b\leq 2a$, note that
@@
$\sum_{\frac{T}{\log^{2}T}<n<T}n^{it}=\Big{(}\sum_{\frac{T}{\log^{2}T}<n\leq\frac{2T}{\log^{2}T}}+\sum_{\frac{2T}{\log^{2}T}<n\leq\frac{4T}{\log^{2}T}}+\cdots+\sum_{T/2<n<T}\Big{)}n^{it}\ll T^{1/2}\log^{3}T.$ (5)
By partial summation, it follows from (5) that
$\sum_{\frac{T}{\log^{2}T}<n<T}n^{\sigma-1+it}\ll T^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{3}T$ (6)
@@
for $\sigma\in\mathbb{R},|\sigma|\leq A$ where $A>0$ is a constant. Combining (1) and (6) yields $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{4}\ll T^{1/2}.$ By the log-derivative of the functional equation for $\zeta$ (see e.g [3, p.403]), note that
$\frac{\zeta^{\prime}}{\zeta}(-\sigma+it)\ll\log t$ (7)
for $\sigma\geq\varepsilon.$ Combining (6) and (7) gives $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{3}\ll T^{1/2}\log^{4}T.$ Thus we arrive at
$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{4}T.$ (8)
For $n\geq m+1$, note that $\frac{1}{\log(n/m)}=\frac{1}{\log n\Big{(}1-\frac{\log m}{\log n}\Big{)}}=\frac{1}{\log n}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log n}\Big{)}^{k}.$ Thus we have
$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=\sum_{k=0}^{\infty}(\log m)^{k}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}(\log n)^{k+1}}.$ (9)
By partial summation, it follows from (5) that
$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}(\log n)^{k+1}}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}(\log T)^{k+1}}.$ (10)
Inserting (10) into the right-hand side of (9) and invoking the fact that $\frac{1}{\log T}\sum_{k=0}^{\infty}\Big{(}\frac{\log m}{\log T}\Big{)}^{k}=\frac{1}{\log(T/m)}$, gives
$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}\log(T/m)}$ (11)
for $T>n\geq m+1>\frac{T}{\log^{2}T}$. The case $\frac{T}{\log^{2}T}<n\leq m-1<T$ is treated similarly by writing $\frac{1}{\log(n/m)}=-\frac{1}{\log m\Big{(}1-\frac{\log n}{\log m}\Big{)}}=-\frac{1}{\log m}\sum_{k=0}^{\infty}\Big{(}\frac{\log n}{\log m}\Big{)}^{k}$ and proceeding exactly as before, so that the same bound holds. Hence for $n\geq m+1$ and $n\leq m-1$, we have
$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}\log(T/m)}.$ (12)
For $\Re(s)>1$, note that $-\frac{\zeta^{\prime}}{\zeta}(s)=\sum_{m=1}^{\infty}\frac{\Lambda(m)}{m^{s}}$. Putting this into $I_{1}$ gives
$I_{1}=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{e\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{e\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (13)
Summing (13) for $\frac{T}{\log^{2}T}<n<T$ and invoking (12) yields
$\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+O(T^{1/2}\log^{6}T)$ (14)
$=-\frac{T}{\pi}\log\log T+o(T)$ (15)
since $H=2T+O(1)$ and $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant [5, p.182]. By Theorem 1 of [3], we have
$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ (16)
uniformly for $2\leq n\ll T$. Define $N(y):=\sum_{0<\gamma\leq y}1$. Note that $N(y)=(y/2\pi)\log(y/2\pi)-y/2\pi+O(\log y)$ [5, Corollary 14.3]. Thus, summing (16) for $1\leq n\leq\frac{T}{\log^{2}T}$ gives
$\sum_{1\leq n\leq\frac{T}{\log^{2}T}}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=N(H)-N(T)-\frac{T}{2\pi}\sum_{2\leq n\leq\frac{T}{\log^{2}T}}\frac{\Lambda(n)}{n}+o(T)$ (17)
$=\frac{T}{\pi}\log\log T+\frac{T}{2\pi}(B-1+\log 2)+o(T).$ (18)
Combining (18), (15), (8) and (3) yields
$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{2}T}}+\sum_{\frac{T}{\log^{2}T}<n\leq T}\Big{)}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ (19)
$\gg T.$ (20)
@@
$\sum_{1\leq n\leq T}n^{\rho_{T}-1}=\zeta(1-\rho_{T})+\frac{T^{\rho_{T}}}{\rho_{T}}+O(T^{\Theta-1})\ll T^{\Theta-1}$ (21)
uniformly for $T\leq\gamma_{T}\leq H.$ Thus
$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}$ (22)
$\ll T^{\Theta}\log T.$ (23)
But (23) is a contradiction to (20), thus our supposition must be false. This completes the proof. ∎
@@
Remark 1. Notice that our argument actually shows that, for all large enough $T\in\mathbb{R}^{+}$, there exists some $\rho_{T}$ with $T\leq\gamma_{T}\leq 2T+1$ and $\Re(\rho_{T})=1-o(1)$.
@@
(i) The functional equation for $\zeta$, and
(ii) Theorem 4.11 of [6]
Both (i) and (ii) do not hold for Beurling zeta functions corresponding to Beurling number systems whose integers are not the ordinary integers. Hence our approach wouldn’t work for such, whose analogues of the RH are known to be true. Also, (ii) does not hold for the Weil zeta functions, which are known to satisfy the RH.